OpenStudy (anonymous):

Factor completely 5c5 + 60c4 + 180c3. 5c3(c + 6)(c - 6) 5c3(c2 + 12c + 36) 5c3(c + 6)2 5(c + 6)2

OpenStudy (solomonzelman):

\(\large\color{black}{ 5c^5 + 60c^4 + 180c^3 }\)

OpenStudy (solomonzelman):

In each of the terms, you have a common multiple/factor of: \(\large\color{black}{ 5c^{~3} }\) . Try to factor out of it, for me please.

OpenStudy (anonymous):

okay ill try

OpenStudy (solomonzelman):

take your time if need to.

OpenStudy (anonymous):

\[5c5+60c4+180c3\]

OpenStudy (anonymous):

\[c3\left( 5c2+12c +36 \right)\]

OpenStudy (anonymous):

did i do that correctly?

OpenStudy (anonymous):

i feel like i messed up with the 5c2 and the c3

OpenStudy (solomonzelman):

you mean, \(\large\color{black}{ 5c^{~3}(c^{~2} + 12c+ 36) }\)

OpenStudy (anonymous):

yes thats what i was going to put but i didnt know i should have gone with my gut feeling

OpenStudy (anonymous):

okay so let me break it down again

OpenStudy (anonymous):

i can divide 12 and 36 by 6 to factor is completely right or am i doing it wrong again?

OpenStudy (solomonzelman):

yes, factor the inside.

OpenStudy (solomonzelman):

Factor: \(\large\color{black}{ c^{~2} + 12c+ 36 }\)

OpenStudy (anonymous):

yes!!!! sorry im just happy im getting it lol

OpenStudy (solomonzelman):

you know that 6+6=12, and 6*6=36.

OpenStudy (anonymous):

\[c2+6c +6\]

OpenStudy (anonymous):

right?

OpenStudy (solomonzelman):

\(\large\color{black}{ c^{~2} + 12c+ 36 = }\) \(\large\color{black}{ c^{~2} + 6c+6c+ 36 =c(c+6)+6(c+6)=? }\)

OpenStudy (anonymous):

ohhhh

OpenStudy (solomonzelman):

yes, so \(\large\color{black}{ c^{~2} + 12c+ 36 }\) is factored into?

OpenStudy (anonymous):

a

OpenStudy (solomonzelman):

how do you have any negative factors in \(\large\color{black}{ c^{~2} + 12c+ 36 }\) , if all the terms of this trinomial are positive?

OpenStudy (solomonzelman):

If I had: b(b+5)+5(b+5) , I would combine this to: (b+5)(b+5)

OpenStudy (anonymous):

c idk i'm close to giving up

OpenStudy (solomonzelman):

C is correct actually

OpenStudy (solomonzelman):

\(\large\color{black}{ 5c^{~3}(c^{~2} + 12c+ 36) }\) \(\large\color{black}{ 5c^{~3}(c+6) ^{~2} }\)

OpenStudy (anonymous):

well i was going in between that and a. im sorry for being a little out of it. i just really dont get it. lol thank you though for keeping up with me :)

OpenStudy (solomonzelman):

yw